n a housing project there are 350 households in which English is usually spoken, 50 in which Spanish is usually spoken, and 100 in which the language is other than English or Spanish. A researcher approaches a house at random to conduct an interview. What is the chance that the usual language in that household will NOT be English? (Answer as a decimal number, i.e. 0.15, not 15%)
1
Expert's answer
2020-03-25T13:42:39-0400

"P(NOT English)=\frac{50+100}{350+50+100}=\frac{150}{500}=0.3"